FTIR SPECTROMETERS UTILIZING MID-INFRARED QUANTUM CASCADE LASERS.

CHRISTIAN PFLUGL, LORAN DIEHL, FEDERICO CAPASSO, School of Engineering and Applied Sciences, Harvard University, Cambridge, Massachusetts, 02138; MARK F. WITINSKI, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, Massachusetts, 02138; \underline{SERGEY SHILOV}, PANG WANG, THOMAS TAGUE, Bruker Optics, 19 Fortune Drive, Billerica, Massachusetts, 01821.

A Fabry-Perot Quantum Cascade Laser source can be used in conjunction with a Fourier Transform Infrared Spectrometer to perform spectroscopic experiments that require orders of magnitude more photons than are emitted by a thermally radiant blackbody source. Three proof-of-concept experiments including laser transmission through liquids, transmission through gases over long distances, and reflection from powders and tablets demonstrated how the increased brightness of a Quantum Cascade Laser enables many important avenues in gas and condensed phase analysis.